# Detecting and Interpreting NSFW Prompts in Text-to-Image Models through Uncovering Harmful Semantics

### v2

@@ Abstract
As text-to-image (T2I) models advance and gain widespread adoption, their associated safety concerns are becoming increasingly critical. Malicious users exploit these models to generate Not-Safe-for-Work (NSFW) images using harmful or adversarial prompts, underscoring the need for effective safeguards to ensure the integrity and compliance of model outputs. However, existing detection methods often exhibit low accuracy and inefficiency. In this paper, we propose HiddenGuard, an interpretable defense framework leveraging the hidden states of T2I models to detect NSFW prompts. HiddenGuard extracts NSFW features from the hidden states of the model’s text encoder, utilizing the separable nature of these features to detect NSFW prompts. The detection process is efficient, requiring minimal inference time. HiddenGuard also offers real-time interpretation of results and supports optimization through data augmentation techniques. Our extensive experiments show that HiddenGuard significantly outperforms both commercial and open-source moderation tools, achieving over 95% accuracy across all datasets and greatly improves computational efficiency.
1 Introduction
@@
Recent advancements in text-to-image (T2I) models, such as Stable Diffusion (Rombach et al., 2022) and DALL·E 3 (Betker et al., 2023), have demonstrated remarkable capabilities in generating high-quality images. However, the widespread use of these models raises ethical concerns, particularly in the generation of Not-Safe-for-Work (NSFW) content, including sexual, violent, hateful, and other harmful images. Recent studies (Qu et al., 2023; Wang et al., 2024; Rando et al., 2022) reveal that users can easily produce NSFW images using malicious prompts, known as NSFW prompts. Consequently, effectively defending against NSFW prompts becomes a crucial challenge.
@@
Among various existing defense methods, input detection is a prominent defense strategy that preemptively identifies unsafe prompts to avoid resource-intensive image generation. Due to its efficiency, this approach has been widely adopted in many text-to-image systems. Existing input detection methods fall into two categories: prompt-based and embedding-based detection. Prompt-based detection (OpenAI, 2023) directly examines the input but often relies on target-model-agnostic classifiers, leading to misclassification due to poor alignment with specific T2I models. Embedding-based detection (Yang et al., 2024b; Liu et al., 2024) leverages the text encoder’s embeddings for better alignment but struggles to capture the deeper semantic information encoded in the text encoder’s attention mechanisms. These limitations suggest that both raw text and final embedding vectors are insufficient for robustly distinguishing benign from malicious prompts, especially against adversarial attacks. This motivates a deeper inspection of the model’s internal representations.
@@
To address the issues above, this paper introduces HiddenGuard, a defense framework designed to detect and interpret NSFW prompts by leveraging the intermediate hidden states of the T2I model. Following current adversarial attacks on T2I models (Yang et al., 2024c; a), our defense focuses on the text encoder. Our key insight is that while NSFW semantics may be entangled and difficult to separate in the final embedding space, they form linearly separable clusters within the hidden state representations of specific layers and attention heads. Building on this discovery, HiddenGuard identifies directional features within these hidden states that robustly encapsulate NSFW semantics. By assessing the magnitude of the input prompt’s hidden state components along these NSFW features, we can effectively detect potential NSFW prompts.
@@
To verify whether NSFW features genuinely represent NSFW semantics and help users understand why a prompt triggers the defense mechanism, we develop a novel interpretability framework based on detection. Our approach considers both textual and image perspectives. On the textual side, we identify NSFW tokens within prompts by leveraging the NSFW features in each attention head. On the image side, we iteratively remove harmful semantics from the hidden states to produce relatively benign embeddings. By generating images from these modified embeddings, we can observe the progressive eradication of NSFW semantics.
@@
Experiment results indicate that HiddenGuard exhibits strong defense capabilities across various text encoders in different T2I models. It surpasses four commercial models, two open-source models, and two state-of-the-art (SOTA) models designed for NSFW prompt detection in both effectiveness and efficiency. Furthermore, HiddenGuard achieves excellent results with minimal data for training and optimization, and it effectively defends against unknown adversarial and adaptive attacks. Additionally, our interpretative approach accurately identifies NSFW tokens and effectively removes NSFW semantics while preserving benign information within the embeddings.
@@
(2) We leverage NSFW features for NSFW prompts detection within T2I models, demonstrating high effectiveness, strong adaptability, excellent optimization ability, and superior efficiency.
@@
(4) We integrate the aforementioned techniques into a unified framework and conduct extensive experiments. The results demonstrate that HiddenGuard outperforms four commercial, two open-source, and two SOTA models.
2 Related Work
@@
Adversarial attacks against T2I models. Adversarial attack methods against T2I models are mainly divided into white-box and black-box approaches. White-box methods primarily utilize the model’s text encoder to optimize the prompt, ensuring that the generated prompt semantically aligns closely with a target prompt containing explicit NSFW information (Yang et al., 2024a). Black-box methods perturb the prompt to find alternative tokens that can replace sensitive words (Yang et al., 2024c; Ba et al., 2023; Deng & Chen, 2023). These methods often utilize reinforcement learning or assistance from large language models to accelerate the search process. Other attack strategies target T2I models with removed concepts (Tsai et al., 2024; Chin et al., 2024). These strategies demonstrate that T2I models can still generate NSFW images even after removing NSFW concepts.
@@
Defensive methods against attacks. Existing defense methods against attacks in T2I models fall into two categories: internal and external safeguards (Zhang et al., 2024). Internal safeguards aim to disable the model’s ability to generate NSFW images by fine-tuning the model itself. Model editing methods (Li et al., 2024b; Poppi et al., 2024; Heng & Soh, 2024; Wu et al., 2024; Qiu et al., 2024) aim to modify the internal parameters by training. However, these methods typically require prolonged training periods and can impact the quality of the generated images. Inference guidance methods (Schramowski et al., 2023; Li et al., 2024a) focus on modifying internal features during the inference stage. They are tuning-free and plug-in, which can be easily inserted into any model. However, the performance of this approach are less robust than model editing.
@@
External safeguards aim to detect potential malicious samples, focusing on input prompts or output images. Output detection (CompVis, 2023; Qu et al., 2023; Schramowski et al., 2022) entails reviewing the generated images to identify NSFW samples. They incur significant inference costs since images must be generated before assessment. Input detection can be classified into prompt-based and embedding-based detection. Prompt-based detection (michellejieli, 2022; unitaryai, 2022) screens prompts to identify those likely to generate NSFW images. They are widely used by online services (Midjourney, 2025; Leonardo.Ai, 2025). Embedding-based detection (Yang et al., 2024b; Liu et al., 2024) examines the conditional embeddings to filter out malicious samples. However, they are highly susceptible to adversarial attacks. In this paper, we primarily focus on input detection within external safeguards, aiming to identify potential NSFW prompts before image generation.
3 Design Intuition
@@
Figure 2: Example 1 corresponds to the prompt “A woman stands on the beach, facing the sea.” Example 2 corresponds to the prompt “A naked woman stands on the beach, facing the sea.”
In this section, we explore how NSFW semantics are concealed within the conditional embeddings and revealed in the text encoder’s hidden states. Our investigation focuses on the CLIP model, the most widely used text encoder in T2I models and the primary target of adversarial attacks. We classify the T2I model prompts into three categories: benign prompts, which do not generate NSFW images; regular NSFW prompts, which are manually crafted with explicit NSFW semantics; and adversarial prompts, which are generated by adversarial attacks and are challenging for humans to interpret as NSFW. We compile a dataset containing over 35,000 prompts across all three categories. A Principal Component Analysis (PCA) of the pooled embeddings of these prompts reveals that while the distributions for benign and NSFW prompts differ, significant overlap prevents reliable differentiation. As shown in 2, two semantically similar prompts can have vastly different safety implications, with one being benign while the other is used to generate NSFW content.
@@
To more accurately delineate the boundary between NSFW and benign prompts, we examine the internal workings of the model by exploring the hidden states across different layers and attention heads. As shown in Figure 1, the first column illustrates the overall distribution of different categories of prompts. While data distributions in some heads overlap significantly, others establish a clear decision boundary, effectively concentrating on NSFW semantics to separate the prompt types.
Furthermore, we observe that individual attention heads exhibit specialized sensitivities to distinct NSFW categories like violence, hate, or sexual content. This suggests a head may excel at identifying a specific subtype even if it struggles to separate NSFW prompts generally. Based on these findings, we propose a method that aggregates signals across all attention heads to achieve a more robust and accurate classification of prompts.
4 Method
Drawing from the insights in Section 3, we develop a framework for detecting and interpreting NSFW prompts. The overall architecture is illustrated in Figure 3. The framework consists of three primary components: the extraction of NSFW features, the detection of NSFW prompts, and the interpretation of the detection process and results.
@@ 4 Method
Figure 3: The overall framework of HiddenGuard. The blue arrows represent the training process, where HiddenGuard extracts NSFW features from the hidden states. The green arrows indicate the inference process, where the prompt passes through the text encoder for detection and interpretation; if it passes, image generation proceeds, otherwise, generation is denied and an interpretation is provided. The red arrows indicate the data augmentation process, where NSFW prompts that successfully bypass detection are added to the training data for data augmentation.
4.1 NSFW Features
@@
In this framework, the ATT layer employs $H$ attention heads to extract information and integrates them into a vector through linear projection. The MLP layer further refines them to obtain intermediate embeddings that represent the overall information of the prompt. However, some information may be obscured or discarded during this process. Therefore, we need to utilize the original outputs $\text{ATT}^{l}(Z^{l-1})$ to extract NSFW semantics effectively.
Considering that $\text{ATT}^{l}(Z^{l-1})$ is a composite of outputs from multiple attention heads, we need to decompose its computational process to extract the information each head represents. Since the CLIP model’s self-attention block employs a causal mask (Radford et al., 2021), only the EOS token holds the complete semantics. Therefore, we concentrate solely on the EOS token. We formalize the ATT output as a sum over $H$ independent attention heads and $N+1$ tokens:
@@
$x_{i,EOS}^{l,h}$ $=\alpha_{i,EOS}^{l,h}W^{l,h}z_{i}^{l-1}.$
where $W^{l,h}$ are transition matrices and $\alpha_{i,EOS}^{l,h}$ are attention weights from the $i$-th token to the EOS token. In this way, we can get the contribution of $h$-th head in $l$-th layer for $h\in[1,H],l\in[1,L]$, which can be expressed as $c^{l,h}=\sum_{i=0}^{N}x_{i,EOS}^{l,h}$.
@@
Table 1: Text-based interpretation of regular NSFW prompts and adversarial prompts. We highlight the words with the strongest NSFW connotations in red, followed by orange and yellow.
To calculate NSFW features, we introduce two sets of prompts: benign prompts and NSFW prompts, where NSFW prompts include regular NSFW prompts and adversarial prompts. We input these prompts into the CLIP model and obtain the output of each attention headOur objective is to maximize the similarity between the NSFW feature and NSFW prompts, while simultaneously minimizing its similarity to benign prompts. We employ the Linear Discriminant Analysis (LDA) algorithm to derive the NSFW feature. The detailed solution process is provided in the Appendix A.2.
4.2 NSFW Prompts Detection
By utilizing the identified NSFW features, we can detect NSFW prompts. We consider the intermediate embeddings of prompts from each attention head as a linear combination of concepts. The projection of these embeddings onto each concept direction represents the contribution of that concept to the embedding. Based on this, we define the projection of the embedding onto the NSFW feature as the NSFW score of a prompt $p$:

### v3

# AEIOU: A Unified Defense Framework against NSFW Prompts in Text-to-Image Models

@@ Abstract
As text-to-image (T2I) models advance and gain widespread adoption, their associated safety concerns are becoming increasingly critical. Malicious users exploit these models to generate Not-Safe-for-Work (NSFW) images using harmful or adversarial prompts, underscoring the need for effective safeguards to ensure the integrity and compliance of model outputs. However, existing detection methods often exhibit low accuracy and inefficiency.
In this paper, we propose AEIOU, a defense framework that is Adaptable, Efficient, Interpretable, Optimizable, and Unified against NSFW prompts in T2I models. AEIOU extracts NSFW features from the hidden states of the model’s text encoder, utilizing the separable nature of these features to detect NSFW prompts. The detection process is efficient, requiring minimal inference time. AEIOU also offers real-time interpretation of results and supports optimization through data augmentation techniques. The framework is versatile, accommodating various T2I architectures. Our extensive experiments show that AEIOU significantly outperforms both commercial and open-source moderation tools, achieving over 95% accuracy across all datasets and improving efficiency by at least tenfold. It effectively counters adaptive attacks and excels in few-shot and multi-label scenarios.
Disclaimer: This article includes potentially disturbing Not-Safe-for-Work (NSFW) text and images. All NSFW images are generated by artificial intelligence. We provide these examples to illustrate how harmful prompts can lead T2I models to generate NSFW content. We have masked and blurred these images, but reader discretion is advised.
I Introduction
@@
Recent advancements in text-to-image (T2I) models, such as Stable Diffusion [rombach2022high], DALL·E 3 [betker2023improving], and Flux [flux], have demonstrated remarkable capabilities in generating high-quality images. However, the widespread use of these models raises significant ethical concerns, particularly in the generation of Not-Safe-for-Work (NSFW) content, including sexual, violent, hateful, and other harmful images. Recent studies [qu2023unsafe, moderator, rando2022red] reveal that users can easily produce NSFW images using malicious prompts, known as NSFW prompts. Consequently, effectively defending against NSFW prompts becomes a crucial challenge.
Existing defense methods can be categorized into internal safeguards, which modify the model to diminish its ability to generate NSFW images, and external safeguards, which detect NSFW content [zhang2024adversarial]. Given that modifying the model may significantly impact the quality of generated content [yang2024guardt2i, gandikota2023erasing], most companies [betker2023improving, Midjourney] tend to adopt detection-based methods, including output and input detection. Output detection [safetysd, qu2023unsafe, schramowski2022can] analyzes the generated NSFW images. However, this approach requires completing the entire generation process, leading to significant resource consumption. Input detection can identify potential NSFW prompts before the generation process, thereby avoiding excessive resource consumption. This is also the focus of our research in this paper.
@@
Existing input detection methods mainly fall into two categories: prompt-based and embedding-based detection. Prompt-based detection [openai, Azure, AWS, Ali] directly examines the input but often relies on target-model-agnostic classifiers, leading to misclassification due to poor alignment with specific T2I models. Embedding-based detection [yang2024guardt2i, liu2024latent] leverages the text encoder’s embeddings for better alignment but struggles to capture the deeper semantic information encoded in the text encoder’s attention mechanisms.
Furthermore, a set of overarching challenges continues to impede the advancement of both prompt-based and embedding-based techniques, impacting their real-world applicability. The first of these is a pronounced vulnerability to adversarial attacks, which can manipulate model outputs through subtle input perturbations, thereby fundamentally compromising the system’s reliability and security [yang2024mma, yang2024sneakyprompt]. A second major issue is architectural over-specialization. The predominant focus on the Stable Diffusion v1 model [sd1.4] means that many proposed methods lack proven applicability across the diverse and rapidly evolving ecosystem of generative architectures, severely limiting their practical utility and scalability. Finally, these methods are characterized by a significant deficit in interpretability. The internal mechanisms governing both content generation and detection are largely opaque, creating a formidable barrier to user trust, accountability, and the diagnosis of failure modes.
@@
To address the issues above, this paper introduces AEIOU, an Adaptable, Efficient, Interpretable, Optimizable and Unified defense framework against NSFW prompts in T2I models. Since current adversarial attacks on T2I models primarily target the text encoder [yang2024sneakyprompt, yang2024mma], as a countermeasure, our defense framework also focuses on the text encoder. Specifically, we analyze the distribution of the text encoder’s hidden states within the feature space, revealing that NSFW prompts present explicit NSFW semantics across various layers and attention heads. Previous research [gandelsmaninterpreting] indicates that hidden states of text encoders can be separated into embeddings for different concepts. Building on this, we analyze the general features of NSFW prompts to identify directions within various attention heads that encapsulate NSFW semantics, termed the NSFW features of the attention heads. By assessing the magnitude of the input prompt’s hidden state components along these NSFW features, we can effectively detect potential NSFW prompts.
@@
To verify whether NSFW features genuinely represent NSFW semantics and help users understand why a prompt triggers the defense mechanism, we develop an interpretability framework based on detection. Our approach considers both textual and image perspectives. On the textual side, we identify NSFW tokens within prompts by leveraging the NSFW features in each attention head. On the image side, we iteratively remove harmful semantics from the hidden states to produce relatively benign embeddings. By generating images from these modified embeddings, we can observe the progressive eradication of NSFW semantics.
Overall, AEIOU overcomes the deficiencies of previous defense methods, demonstrating not only high accuracy but also exhibiting the following five characteristics:
(1) Adaptable. By analyzing attention heads within the text encoder to specifically capture NSFW features, AEIOU can adapt to any transformer-based text encoder, allowing application across various T2I model architectures.
(2) Efficient. AEIOU operates without complex models for detection, requiring minimal time for both training and inference, ensuring high efficiency.
(3) Interpretable. The inference process of AEIOU does not involve black-box models; instead, it relies on comparisons with NSFW features. The results can be visualized and interpreted across text and image modalities, offering both process and outcome interpretability.
(4) Optimizable. By assigning greater weight to red-teaming data, we can achieve data augmentation for AEIOU with only a few samples, simplifying its updating process.
(5) Unified. AEIOU integrates training, inference, interpretation, and further optimization into a unified framework, avoiding isolated processes.
@@
Experiment results indicate that AEIOU exhibits strong defense capabilities across various text encoders in different T2I models. It surpasses four commercial models, two open-source models, and two state-of-the-art models designed for NSFW prompt detection in both effectiveness and efficiency. Furthermore, AEIOU achieves excellent results with minimal data for training and optimization, and it effectively defends against unknown adversarial and adaptive attacks. Additionally, our interpretative approach accurately identifies NSFW tokens and effectively removes NSFW semantics while preserving benign information within the embeddings.
@@
(2) Based on the insights above, we leverage NSFW features for NSFW prompts detection within T2I models, demonstrating high effectiveness, strong adaptability, excellent optimization ability, and superior efficiency.
@@
(4) We integrate the aforementioned techniques into a unified framework and conduct extensive experiments. The results demonstrate that AEIOU outperforms four commercial, two open-source, and two state-of-the-art models.
II Related Work
II-A Adversarial Attacks against T2I Models
With the continuous advancement of T2I models, they are becoming increasingly integrated into various aspects of daily life and work [ho2020denoising, rombach2022high, betker2023improving, podell2023sdxl, ruiz2023dreambooth, esser2024scaling, liu2023riatig]. Despite their strengths, they are susceptible to adversarial attacks that modify prompts to sneak past defenses and produce NSFW content like pornography, violence or politically sensitive imagery [tsai2023ring, chin2023prompting4debugging, deng2023divide].
@@ II-A Adversarial Attacks against T2I Models
Present adversarial attack methods are mainly divided into white-box and black-box approaches. White-box methods primarily utilize the model’s text encoder to optimize the prompt, ensuring that the generated prompt semantically aligns closely with a target prompt containing explicit NSFW information, even without sensitive words [yang2024mma]. In contrast, black-box methods perturb the prompt to find alternative tokens that can replace sensitive words [yang2024sneakyprompt, ba2023surrogateprompt]. These methods often utilize reinforcement learning or assistance from large language models to accelerate the search process. In addition, some attack strategies target T2I models with removed concepts [tsai2023ring, chin2023prompting4debugging]. These strategies demonstrate that T2I models can still generate NSFW images even after removing NSFW concepts.
Current adversarial methods have proven highly effective. Consequently, it is crucial to develop robust defense mechanisms to counter these attack strategies and ensure the safe and responsible use of T2I models.
II-B Defensive Methods against Attacks
Based on the knowledge of the model, existing defense methods against attacks in T2I models fall into two categories: internal and external safeguards [zhang2024adversarial].
@@ II-B Defensive Methods against Attacks
Internal safeguards aim to disable the model’s ability to generate NSFW images by fine-tuning the model itself. They can be divided into model editing and inference guidance. Model editing methods [gandikota2023erasing, kumari2023ablating, li2024safegen, poppi2024safe, heng2024selective, orgad2023editing, wu2024unlearning, hu2025safetext] aim to modify the internal parameters by training. However, these methods typically require prolonged training periods, and parameter modifications can impact the quality of the generated images. Moreover, most methods focus solely on safety against malicious prompts while ignoring the adversarial prompts. In contrast, inference guidance methods [schramowski2023safe, li2024self] focus on modifying internal features during the inference stage. Unlike model editing methods, they are tuning-free and plug-in, which can be easily inserted into any model. However, these methods also fail to account for adversarial prompts, making them susceptible to targeted attacks.
@@
External safeguards aim to filter out potential malicious samples by examining intermediate variables during the generation process. Current detection methods focus on prompts, conditional embeddings, or generated images. Image-based moderation [safetysd, qu2023unsafe, schramowski2022can] entails reviewing the generated images to identify NSFW samples. They incur significant inference costs since images must be generated before assessment. Prompt-based moderation [openai, AWS, Azure, NSFWtext, Detoxify] screens input prompts to identify those likely to generate NSFW images. Given their lower cost, they are widely used by online services like Midjourney [Midjourney] and Leonardo.Ai [Leon]. Nonetheless, these methods generally lack targeted defenses against adversarial attacks, making them susceptible to circumvention. Embedding-based moderation [yang2024guardt2i, liu2024latent] examines the conditional embeddings to filter out malicious samples. While this approach offers some resistance to adversarial attacks, it relies on large-scale models for classification, resulting in high costs. Additionally, it also suffers from low accuracy and remains vulnerable to adaptive attacks.
II-C Interpretation on the CLIP Model
As the most widely used text encoder in T2I models, the CLIP model [radford2021learning] is a significant focus of adversarial attack research. Recent studies [gandelsmaninterpreting, bhalla2024interpreting, zhao2024gradient, aflalo2022vl] have investigated the CLIP model to analyze its internal mechanisms. For instance, Bhalla et al. [bhalla2024interpreting] found that the embeddings generated by CLIP exhibit strong linear properties and can be decomposed into combinations of various concepts. Gandelsman et al. [gandelsmaninterpreting] discovered that different attention heads within the CLIP model are responsible for interpreting different semantics, which are then combined to produce the final output embeddings. However, these studies primarily focus on the image domain. In this paper, we build upon existing work to further explore the properties of the CLIP model in the text domain and propose a novel interpretation method to interpret how prompts containing NSFW semantics are generated.
III Method
Figure 1: Left: The distribution of output embeddings from CLIP model. Right: Example 1 corresponds to the prompt “A woman stands on the beach, facing the sea.” Example 2 corresponds to the prompt “A naked woman stands on the beach, facing the sea.” On the right are the images generated from them.
III-A Design Intuition
In this section, we explore how NSFW semantics are concealed within the conditional embeddings and revealed in the text encoder’s hidden states. Our investigation focuses on the CLIP model [radford2021learning], the most widely used text encoder in T2I models and the primary target of adversarial attacks.
@@ III-A Design Intuition
We classify the T2I model prompts into three categories: benign prompts, which do not generate NSFW images; regular NSFW prompts, which are manually crafted with explicit NSFW semantics; and adversarial prompts, which are generated by adversarial attacks and are challenging for humans to interpret as NSFW. Through extensive data collection, we compiled a dataset containing over 35,000 prompts across all three categories. We utilize the CLIP model to obtain pooled embeddings of these prompts and examine their distribution with a PCA map [abdi2010principal], as depicted in Figure 1. Although the distributions of benign and NSFW data differ, considerable overlap makes effective differentiation challenging.
This phenomenon is primarily attributed to the architecture of the CLIP network, which comprises multiple layers and attention heads. Each layer employs a multi-head self-attention mechanism [vaswani2017attention], where each head independently extracts information from the prompt. The outputs of these heads are then combined through a linear layer and passed to the next layer. Linear layers integrate information from various heads, making it challenging to isolate individual pieces of information. As the network’s depth increases, the outputs from different layers are accumulated through residual connections, exacerbating the entanglement of information. Consequently, benign prompts and NSFW prompts become intermixed within the pooled embedding space. As illustrated in the figure, two sentences describing the same object might appear semantically similar; however, one could be used to generate NSFW images while the other remains benign.
@@
To more accurately delineate the boundary between NSFW and benign prompts, we examine the internal workings of the model by exploring the hidden states across different layers and attention heads. Figure 2 illustrates the distribution of benign, regular NSFW and adversarial data across several attention heads. The first column illustrates their overall distribution. In some heads, significant overlap remains between benign and NSFW data distributions. However, in other heads, a distinct boundary between these data types is evident. This reveals that attention heads within the model exhibit differing sensitivities to NSFW content. Certain attention heads concentrate on the NSFW semantics within prompts, allowing them to differentiate between benign and NSFW prompts effectively.
We also find that different NSFW prompt categories are processed uniquely across attention heads. Columns two, three, and four of Figure 2 illustrate how violence, hate, and sex categories differ from benign data. In the first attention head, violence and hate data are separable from benign data, whereas sex data shows overlap. The second attention head clearly differentiates violence from benign data, with weaker distinctions in the other categories. In the third head, sex and hate data are distinctly separated from benign data, while violence data overlaps significantly. These findings suggest that attention heads specialize in handling specific NSFW content types. Even if a head struggles to differentiate between benign and NSFW data, it may still excel at distinguishing a particular type of NSFW prompt.
Based on these findings, we propose leveraging the attention heads within the text encoder to differentiate between benign and NSFW prompts effectively. By capitalizing on the diverse focus of different attention heads, we can aggregate information from all attention heads to achieve more accurate prompt classification.
III-B Framework Overview
Based on the findings from Section III-A, we develop a framework for detecting and interpreting NSFW prompts. The overall architecture of this framework is illustrated in Figure 3. Initially, we identify the direction within each attention head’s hidden states that best represent NSFW semantics by analyzing the distribution differences between benign and NSFW prompts, which we call NSFW features. When a new prompt is inputted, we can assess the risk of generating NSFW images by evaluating the projection magnitude of the prompt along the NSFW feature. To interpret our assessment, we pinpoint the token most representative of NSFW semantics within the prompt, analyzing why the current prompt is classified as an NSFW prompt. Furthermore, we can progressively eliminate NSFW semantics from the hidden states and input the modified prompt embeddings to generate images. This process enables us to observe how NSFW semantics are gradually removed from the embeddings. Finally, we can conduct red team testing on the framework or monitor it in real-time post-deployment to collect prompts that bypass current defenses. By incorporating them into the training set, we achieve more accurate NSFW feature identification and detection results.
The practical application of this framework manifests in two main ways. First, it prevents NSFW image generation from the outset when malicious users attempt to employ adversarial prompts. Second, when regular users inadvertently input prompts with NSFW semantics and encounter blockages, we can interpret why the prompt is considered inappropriate. This helps users quickly identify problematic tokens and modify their prompts to generate the desired image.
@@ III-B Framework Overview
Figure 3: The overall framework of AEIOU. The blue arrows represent the training process, where data enters the text encoder, and AEIOU extracts NSFW features from the hidden states. The green arrows indicate the inference process, where the prompt passes through the text encoder for detection and interpretation; if it passes, image generation proceeds, otherwise, generation is denied and an interpretation is provided. The red arrows indicate the data augmentation process, involving red-teaming tests on AEIOU, where NSFW prompts that successfully bypass detection are added to the training data for data augmentation.
III-C NSFW Features
@@
In this framework, the ATT layer employs $H$ attention heads to extract information and integrates them into a vector through linear projection. The MLP layer further refines them to obtain intermediate embeddings that represent the overall information of the prompt. However, some information may be obscured or discarded during this process. Therefore, we need to utilize the original outputs from the attention heads to extract NSFW semantics effectively.
Considering that the outputs from multiple attention heads have been preliminarily integrated into $\text{ATT}^{l}(Z^{l-1})$, we need to decompose its computational process to extract the information each head represents. Since the CLIP model’s self-attention block employs a causal mask [radford2021learning], only the EOS token holds the complete semantics. Therefore, we concentrate solely on the EOS token. Following Elhage et al. [elhage2021mathematical] and Gandelsman et al. [gandelsmaninterpreting], we formalize the ATT output as a sum over $H$ independent attention heads and $N+1$ tokens:
@@
$x_{i,EOS}^{l,h}$ $=\alpha_{i,EOS}^{l,h}W_{VO}^{l,h}z_{i}^{l-1}.$
where $W_{VO}^{l,h}$ are transition matrices and $\alpha_{i,EOS}^{l,h}$ are attention weights from the $i$-th token to the EOS token. In this way, we can get the contribution of $h$-th head in $l$-th layer for $h\in[1,H],l\in[1,L]$, which can be expressed as $c^{l,h}=\sum_{i=0}^{N}x_{i,EOS}^{l,h}$.
@@
TABLE I: Text-based interpretation of regular NSFW prompts and adversarial prompts with CLIP-ViT-L. We highlight the words with the strongest NSFW connotations in red, followed by orange and yellow.
@@
where $\mu_{b}^{l,h}$ and $\mu_{m}^{l,h}$ are the mean value of $\{c_{b_{k}}^{l,h}\}$ and $\{c_{m_{k}}^{l,h}\}$. In summary, our goal is to maximize the distance between the projected means of $\{c_{b_{k}}^{l,h}\}$ and $\{c_{m_{k}}^{l,h}\}$ on the vector $u^{l,h}$, while simultaneously minimizing their respective variances. We can employ Linear Discriminant Analysis (LDA) [balakrishnama1998linear] to solve this problem, ultimately obtaining the NSFW feature $\{u^{l,h}\}$:
$u^{l,h}=S_{w}^{-1}(\mu_{m}^{l,h}-\mu_{b}^{l,h}),$ (6)
$S_{w}=\sum_{k=0}^{K_{m}}\|c_{m_{k}}^{l,h}-\mu_{m}^{l,h}\|^{2}+\sum_{k=0}^{K_{b}}\|c_{b_{k}}^{l,h}-\mu_{b}^{l,h}\|^{2},$ (7)
where $K_{b}$ is the number of benign prompts, and $K_{m}$ is the number of malicious prompts. In text encoders other than CLIP, such as T5 [raffel2020exploring], we can still use the method above to extract NSFW features. Any text encoder utilizing multi-head self-attention can be adapted to this approach.
III-D NSFW Prompts Detection
By utilizing the identified NSFW features, we can detect NSFW prompts. According to previous research [gandelsmaninterpreting], we can consider the intermediate embeddings of prompts in the CLIP model as a linear combination of concepts. The projection of these embeddings onto each concept direction represents the contribution of that concept to the embedding. Based on this, we define the projection of the embedding onto the NSFW feature as the NSFW score of a prompt $p$: